OpenStudy (anonymous):

The length of a rectangle is 5.5 feet. If the area is 44 square feet, what is the width of the rectangle, in feet? A. 242 B. 8 C. 16.5 D. 38.5

OpenStudy (anonymous):

@JA1

OpenStudy (ja1):

@melody16 @tcarroll010 @mathstudent55

OpenStudy (ja1):

So first we need the formula for the area of a rectangle.

OpenStudy (anonymous):

Which is what formula?

OpenStudy (ja1):

A = L * W

OpenStudy (ja1):

So let's substitute some numbers here.

OpenStudy (ja1):

What is the length?

OpenStudy (anonymous):

5.5

OpenStudy (ja1):

Perfect so A = 5.5 * W

OpenStudy (ja1):

Now what is the area? (A)

OpenStudy (anonymous):

44.

OpenStudy (ja1):

good so we need to know what number times 5.5 is 44

OpenStudy (ja1):

So let's do the inverse operation... Division

OpenStudy (anonymous):

8

OpenStudy (ja1):

Ok good xD

OpenStudy (anonymous):

8 x 5.5 is 44.

OpenStudy (ja1):

So there you go A =L * W is the same as 44 = 5.5 * 8

OpenStudy (ja1):

So it is B :)

OpenStudy (anonymous):

ohhh, how easy lol, thank you. :)

OpenStudy (ja1):

Do you understand the reasoning?

OpenStudy (ja1):

No problem I am glad to help anytime :) see it wasn't so hard ^,^

OpenStudy (anonymous):

yes, thank you very much. @JA1

OpenStudy (ja1):

No problem :)

OpenStudy (ja1):

BTW did you know you can reward any user who helps you on your question by clicking the "best response" button? Thanks @tcarroll010